OpenStudy (anonymous):

solve the polynomial inequality. Express the solution using interval notation. x(2x+3)<0

OpenStudy (anonymous):

this is a parabola facing up, it will be negative between the zeros. the zeros are 0 and \[\frac{3}{2}\]

OpenStudy (anonymous):

so it will be negative over the interval \[(-\frac{3}{2},0)\]

OpenStudy (anonymous):

oops i made a mistake. i wrote the zeros are 0 and \[\frac{3}{2}\] but is should be \[-\frac{3}{2}\] and 0

OpenStudy (anonymous):

answer is correct though

OpenStudy (anonymous):

thank you.. sry it took so long :)